OpenStudy (anonymous):

abs((1-2x)/x)=3x solve for x

OpenStudy (anonymous):

\[|\frac{1-2x}{x}|=3x\] like that?

OpenStudy (anonymous):

since absolute value is always greater than or equal to zero, we can assume that \(x\) is as well, otherwise there is no solution. therefore you can solve \[\frac{1-2x}{x}=3x\] only in the case where \(x\geq 0\)

OpenStudy (anonymous):

since \(x\neq 0\) and also we know \(x>0\) multiply both sides by \(x\) to get \[1-2x=3x^2\]or \[3x^2+2x-1=0\] and solve the quadratic

OpenStudy (anonymous):

this one actually factors as \((3x-1)(x+1)=0\) so \(x=\frac{1}{3}\) we can throw out the \(-1\) solution because it is negative and \(x\) cannot be negative

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

yw, hope steps are clear